OpenStudy (anonymous):

5. Describe two different methods to estimate 1.35 + 2.4 + 1.48, and give their results. How do these compare to the actual solution?

OpenStudy (anonymous):

can someone please help me i will medal and FAN!!! can some give me the answer

OpenStudy (michele_laino):

please are those experimental measures? and in that case do you have the corresponding uncertainties?

OpenStudy (anonymous):

um i dont know what that means but can you help me get the answer @Michele_Laino

OpenStudy (michele_laino):

I don't know precisely here is my reasoning If I take two decimal figures, for each quantity, I get: \(1.35+2.40+1.48=...\) whereas if I take one decimal figure for each number, then I get: \(1.4+2.4+1.5=...\)

OpenStudy (michele_laino):

as we can see, in the last case I have rounded each number

OpenStudy (anonymous):

okay so is that the answer @Michele_Laino

OpenStudy (michele_laino):

yes! I think so, nevertheless, please ask to another helper

OpenStudy (anonymous):

what are the answer to If I take two decimal figures, for each quantity, I get: 1.35+2.40+1.48=... whereas if I take one decimal figure for each number, then I get: 1.4+2.4+1.5=...

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (jiteshmeghwal9):

well if u round off the inputs u will gt less accurate values

OpenStudy (jiteshmeghwal9):

w/o rounding off the inputs 1.35+2.4+1.48=5.23 rounding the result we gt 5.2

OpenStudy (jiteshmeghwal9):

if we round off the inputs 1.4+2.4+1.5=5.3

OpenStudy (jiteshmeghwal9):

actual value is 5.23 we see the results when we applied these two methods if we estimate the inputs the result will be less inaccurate

OpenStudy (michele_laino):

If I make the computation above, I get this: \[\Large \begin{gathered} 1.35 + 2.4 + 1.48 = 5.23 \hfill \\ \hfill \\ 1.4 + 2.4 + 1.5 = 5.3 \hfill \\ \end{gathered} \] which are the requested results

Directrix (directrix):

@lauren.ascough I was unaware that you were not pleased with our working of this problem an hour or so ago. http://openstudy.com/users/directrix#/updates/56e7b260e4b0c8b420a265ad

OpenStudy (anonymous):

yes i did not understand and you were not answering me so i just asked other people @dirtex

OpenStudy (anonymous):

@directrix

OpenStudy (mathmale):

Lauren: The goal of the problem you have posted is to give you experience with estimating sums. What does the word "estimate" mean to you? Several people have tried, in this and your previous post, to explain that rounding off each of the numbers you're adding together will lead to a reasonably accurate ESTIMATE of the sum of those numbers. Make certain that you understand the concept of estimation first. What does this word mean to you? Rather than ask for answers ("ok so can you send me what i would write for the question"), ask questions so that you can obtain the info you need to solve the problem yourself. OpenStudy is not intended to be a source of straight answers; you are expected to do enough work yourself so as to end up with an understanding of the concepts involved.